OpenStudy (anonymous):

y = 1/2x - 6 2x + 6y = 19 with either elimination substitution or graphing. Thanks :) these three ways of solving are really confusing

OpenStudy (anonymous):

Let us take one at a time..... Substitution This is the easiest of all IF on of the equations is given as y = something. In the above case, it is. So, you take that value of y and substitute it (therefore the word substitution) in the second equation. 2x + 6(1/2*x - 6) = 19 Now, you have an equation in x that can be solved to figure out x. Then, use y = 1/2*x = 6 to figure out value of y.

OpenStudy (anonymous):

? * ?

OpenStudy (anonymous):

Then do i solve 2x + 6 (1/2x -6) = 19? 2x + 3x -6 = 19? 5x -6 = 19? 5x = 13? or is it 5x = 25?

OpenStudy (anonymous):

It is 36 (6 * 6) like this: 2x + 3x -36 = 19?

OpenStudy (anonymous):

oh yeah 6 * 6 ! so then what i was doing before but with the 36 instead?

OpenStudy (anonymous):

Correct. Also, check my reply to that other question about elimination. Hope that helps.

OpenStudy (anonymous):

I would subtract 36 from 19 right ?

OpenStudy (anonymous):

or no add..

OpenStudy (anonymous):

Add because when you have -36 and take it to the other side, you add.

OpenStudy (anonymous):

Okay that makes sense